# Med-GLIP: Advancing Medical Language-Image Pre-training with Large-scale Grounded Dataset

### v1

@@ Abstract
Medical image grounding aims to align natural language phrases with specific regions in medical images, serving as a foundational task for intelligent diagnosis, visual question answering (VQA), and automated report generation(MRG). However, existing research is constrained by limited modality coverage, coarse-grained annotations, and the absence of a unified, generalizable grounding framework. To address these challenges, we construct a large-scale medical grounding dataset Med-GLIP-5M comprising over 5.3 million region-level annotations across seven imaging modalities, covering diverse anatomical structures and pathological findings. The dataset supports both segmentation and grounding tasks with hierarchical region labels, ranging from organ-level boundaries to fine-grained lesions. Based on this foundation, we propose Med-GLIP, a modality-aware grounding framework trained on Med-GLIP-5M. Rather than relying on explicitly designed expert modules, Med-GLIP implicitly acquires hierarchical semantic understanding from diverse training data—enabling it to recognize multi-granularity structures, such as distinguishing lungs from pneumonia lesions. Extensive experiments demonstrate that Med-GLIP consistently outperforms state-of-the-art baselines across multiple grounding benchmarks. Furthermore, integrating its spatial outputs into downstream tasks, including medical VQA and report generation, leads to substantial performance gains. Our dataset will be released soon.
Index Terms:
Medical Dataset, Medical Image Grounding, Grounded Language-Image Pre-training
@@ I Introduction
With the rapid progress of medical AI, establishing precise alignment between natural language descriptions and specific regions in medical images has become a foundational step for tasks such as intelligent diagnosis, surgical navigation, and multimodal question answering [1, 2, 3, 4, 5, 6, 7, 8, 9]. Medical image grounding aims to localize anatomical structures or pathological findings based on language input, enabling spatial-semantic correspondence across modalities. For example, when a radiologist states “a tumor is located in the upper right corner,” the model must accurately identify the corresponding image region. This task not only improves model interpretability but also provides critical spatial priors for downstream applications like visual question answering (VQA) and medical report generation (MRG), playing an essential role in clinical decision support[10, 9, 11, 12, 13, 14, 15, 16]
However, compared to the natural image domain, medical image grounding faces several unique challenges [17, 18]. First, publicly available grounding datasets are extremely scarce, especially those with large-scale, multi-organ, and multi-modality annotations, which significantly limits progress in this field. Second, medical images are highly specialized, and exhibit substantial heterogeneity across modalities such as CT, MRI, and ultrasound—in terms of spatial resolution, anatomical appearance, and contrast—making cross-modal alignment inherently complex [19, 20, 21]. Moreover, many target regions, such as small lesions or vascular branches, often lack clear boundaries and exhibit high anatomical variability, placing greater demands on the model’s precision and generalization capabilities.
@@
CT-RAD 2023 10410 studies 3D-$B$ Volumetric abnormalities
@@
To address data scarcity and semantic misalignment in medical image grounding, prior works have explored both dataset construction and cross-modal modeling. On the dataset side, SLAKE [1] offers region-phrase annotations for chest X-rays, marking an early attempt at grounding but with only 6k samples and limited modality coverage. MedTrinity-25M [22] provides large-scale but loosely aligned image-text pairs across tasks, lacking fine-grained region-level supervision. These datasets commonly suffer from: (1) limited modality and organ diversity; (2) absence of dense region-level annotations; and (3) a focus on classification or QA tasks rather than spatial-semantic alignment. On the modeling side, recent approaches adapt natural-image cross-modal models to the medical domain (e.g., LLaVA-Med [23], MedKLIP [24], MedSAM [25]), often with structured medical knowledge. LLaVA-Med [23] introduces multimodal LLMs but relies on image-caption pairs without fine-grained alignment. MedKLIP [24] employs region-phrase contrastive learning but remains X-ray–specific. MedSAM [25] incorporates structure-aware priors for zero-shot organ segmentation based on SAM [26], yet depends on prompts and lacks deep language grounding. OntoRay further integrates radiological ontologies to capture causal relations between terms and regions [27]. Despite these efforts, a unified, modality-adaptive grounding framework capable of multi-scale semantic alignment and inter-modality generalization is still missing, and its impact on downstream tasks such as VQA or report generation remains underexplored.
To address these challenges, we first propose Med-GLIP-5M, a large-scale medical grounding dataset comprising over 5.3 million region-level annotations across seven imaging modalities, covering diverse anatomical structures and pathological findings. We first construct a comprehensive dataset containing over 5.3 million region-level annotations across seven imaging modalities and dozens of anatomical and pathological categories (Table I). This dataset is curated from multiple public medical segmentation sources and is characterized by three key properties: segmentability, groundability, and multi-granularity. We have standardized and refined the data processing to ensure it adheres to data usage regulations and ethical considerations. Based on this foundation, we also proposed Med-GLIP, which adopts a hierarchical expert paradigm, where dedicated grounding modules are trained for each imaging modality, enabling accurate localization of fine-grained anatomical structures (Figure 4). Furthermore, we demonstrate that incorporating Med-GLIP’s grounding outputs into downstream tasks—such as medical VQA and report generation—leads to significant performance improvements, highlighting the value of explicit spatial grounding for multimodal medical reasoning. In summary, our contributions are as follows:
@@
  • To the best of our knowledge, we introduce the largest and most diverse medical grounding dataset to date, Med-GLIP-5M, spanning 7 modalities, 30+ anatomical regions, and over 5.3 million image-text region pairs, addressing the long-standing data scarcity in this field.
  • We propose Med-GLIP, a modality-aware hierarchical expert framework that enables high-precision grounding across diverse medical imaging types by modeling fine-grained structural differences.
  • Through extensive experiments, we validate the effectiveness of spatial grounding in downstream tasks, showing that Med-GLIP enhances both VQA and report generation performance. Our work provides a unified solution across data, modeling, and application perspectives, advancing the development of generalizable medical vision-language models.
@@ II Med-GLIP-5M Construction
II-A Data Source Analysis
In this study, we curated a comprehensive medical image dataset by aggregating resources from multiple platforms, including publicly available repositories such as GitHub, Grand Challenge, OpenNeuro, SYNAPSE, and BraTS, as well as expert annotations provided by collaborating clinicians from affiliated hospitals. The collected datasets span a wide range of imaging modalities, spatial resolutions, and anatomical regions, ensuring diversity across organs and clinical tasks, as shown in Figure 2. A key principle in our data collection strategy was to include images with multiple segmentation masks wherever possible, enabling multi-label learning and enhanced contextual understanding.
Following data acquisition, we conducted a systematic preprocessing workflow. Datasets were first categorized by organ type and sorted in ascending order of data volume. Duplicated samples and low-quality data—such as images with insignificant mask coverage or corrupted annotations—were filtered out. To resolve label inconsistency, semantically similar class names were standardized into a unified naming scheme. For the remaining clean data, we computed bounding boxes (BBox) from the segmentation masks and enriched each dataset with metadata including label names, modality type, dataset size, and mask-image pairing integrity, as shown in Figure 3 and Figure 4.
@@ II-A Data Source Analysis
Figure 2: Overview of the Med-GLIP-5M dataset construction process. (a) Data sources: We collect 129 public medical datasets from diverse platforms including OpenNeuro, Grand Challenge etc. (b) Data preprocessing pipeline: The datasets are categorized, ambiguities resolved, ROIs and textual labels are added, and metadata is integrated into a unified format. (c) Example visualizations of processed data: Our dataset contains fine-grained ROI annotations and corresponding textual descriptions across multiple modalities, supporting region-level grounding and structured understanding.
II-B Data Collection
We curated a multi-source medical imaging dataset for vision-language grounding. Each image was selected with a preference for multi-instance spatial annotations—either multiple masks or bounding boxes per image—thus ensuring the resulting dataset is well-suited for vision-language grounding tasks. Datasets were extracted from several platforms including but not limited to: GitHub, Grand Challenge , OpenNeuro, BraTS, SYNAPSE, and CodaLab. Selection criteria: (1) DICOM/NIfTI accessibility; (2) spatial annotations (masks/bounding boxes); (3) coverage of multiple anatomical regions. This yielded 2720979 images spanning X-ray, CT, MRI across multiple organs including abdomen(32.8%), bone(23.9%), head(17.3%), chest(7.6%) etc. The final dataset has 4.13 masks/image on average, with various modality including CT(61.84%), MRI(26.71%), Ultrasound(1.44%), X-RAY(0.22%), Dermoscopy(0.21%), Endoscopy(0.09%), Fundus(0.09%).
II-C Data Quality Control
A three-tier quality control (QC) pipeline was implemented post-collection. Firstly, all DICOM images underwent dcm2niix (v1.0) with --check-mosaic flag. Non-readable or broken files were discarded. 3D dataset were sliced into 2D on different axis to improve the available data quantity. We standardized the image format across datasets to PNG, while preserving the original resolution to maintain fidelity with source annotations. Secondly, each image was verified to have a corresponding segmentation mask. We implemented format-specific matching scripts to ensure annotation-visual alignment. Samples lacking a valid associated mask were excluded from further processing. Thirdly, we filtered inconsistent annotations: entries with malformed structure, undefined label semantics, or ambiguous mappings were excluded. Moreover, we removed samples where the annotated mask covered only a negligible portion of the image (empirically set as 1.5% area exclusion according PASCAL VOC min IoU standard), as such cases do not provide meaningful spatial grounding for downstream tasks[28]. This quality assurance workflow ensures that only well-structured, clearly labeled, and spatially meaningful samples are retained for model training and evaluation.
II-D Data Preprocessing
The data preprocessing phase for Med-GLIP-5M was designed to ensure high-quality data while adhering to ethical standards and regulatory requirements. After acquisition, datasets underwent a multi-stage refinement process. Initially, data were categorized by organ type and volume, with duplicates and low-quality samples removed. Semantic label inconsistencies were resolved through standardized naming. Clean data were then augmented with geometric and intensity transformations to enhance diversity without compromising diagnostic integrity. Image formats were standardized to ensure compatibility across machine learning frameworks. A rigorous validation protocol verified the consistency between images and segmentation masks. Additionally, clinical metadata such as patient history and diagnostic outcomes were integrated where available. This structured approach resulted in a dataset that is both large-scale and high-quality, providing a robust foundation for medical AI research.
II-E Data Statistics
The dataset used in this study comprises 198 fine-grained annotation labels, which were analyzed via human and re-organized into 38 broader anatomical categories based on label semantics and corresponding dataset documentation by large language models and verified by human. This process, facilitated by large language models and verified by human experts, aimed to enhance anatomical coherence and interpretability for downstream modeling (e.g., mapping specific hip joints to ’pelvis(hip)’ and optic components to ’head(eye)’). These 38 hierarchical classes span 6 major body regions, with a cumulative image count exceeding 11 million. Notably, the brain, spine, and abdomen represent the systems with the highest annotation volumes. This hierarchical label structure allows for flexible experimentation across tasks such as organ-level segmentation, multi-organ detection, and facilitates domain adaptation and multi-label learning.
The dataset exhibits substantial heterogeneity with 7 distinct imaging modalities. Computed Tomography (CT) and Magnetic Resonance Imaging (MRI) are predominant, contributing approximately 4.2 million and 1.08 million images, respectively. These two modalities collectively account for over 80% of the total dataset volume. CT scans, reported as constituting 78% among modalities in one analysis, provide extensive coverage across multiple organ systems, including the spine, lung, liver, and kidney. MRI, accounting for 20% among images in a similar context, is particularly enriched in high-resolution brain and neuroimaging data, as well as prostate, cardiac, and vascular regions. The remaining modalities, such as ultrasound, contribute important diversity.
@@ III Med-GLIP
In our approach, medical object detection is reconceived as a phrase grounding task, wherein each image region that is identified corresponds to its matching medical phrase. Given a predefined set of medical concepts pertinent to a particular imaging modality, for instance, {“pneumonia”, “nodule”, “fracture”} for X-ray images, a textual prompt is constructed, such as:
@@
Following methodologies similar to GLIP [29], we can employ pre-trained language models, e.g., BERT [30], to encode more semantically rich prompts (e.g., “pneumonia. nodule. fracture.”), which has demonstrated empirical advantages. Within our modality-specific grounding framework, alignment scores $S_{\text{ground}}$ are computed between medical image region features $\mathbf{F}$ and the encoded word or token features $\mathbf{T}$ from the prompt. This is formally expressed as:
@@ IV-A Experiment Setup
Datasets for Region-Level Grounding.  To evaluate the grounding performance of our GLIP-experts, we selected six datasets spanning five imaging modalities: SegRap2023_ct (CT), AMOS2022_MR (MRI), sz_cxr (X-ray), CETUS2014 (Ultrasound), and CVC-ClinicDB (Endoscopy). All datasets except CVC-ClinicDB are sourced from the IMed-361M benchmark[31], which provides standardized train-test splits (90%/10%). For evaluation, we use only the held-out test sets. CVC-ClinicDB lacks predefined splits and is used in its entirety. These datasets collectively cover 5,551 images with rich structural annotations across modalities.
Datasets for MRG.  We adopt two widely-used benchmarks: MIMIC-CXR[32] and IU-Xray[33]. MIMIC-CXR includes 377,110 chest X-ray images and corresponding free-text radiology reports from 227,835 studies. It is evaluated using the MLRG framework to compute both natural language generation (NLG) metrics and concept grounding scores (e.g., CE). IU-Xray, with 7,470 image-report pairs, offers a smaller and more structured alternative, and is evaluated using the R2Gen model on standard NLG metrics.
@@
Datasets for Med-VQA.  We evaluate model generalization to vision-language reasoning via three medical VQA datasets: VQA-RAD[34], SLAKE[1], and PathVQA[35]. These datasets span radiological and pathological domains, include both English and Chinese annotations, and provide over 50,000 QA pairs. During evaluation, GLIP-experts support both: (1) Closed-ended multiple-choice reasoning by verifying spatial alignment between phrases and detected regions; and (2) Open-ended answer generation through grounded semantic reasoning. All evaluations follow official train-test splits for reproducibility and fair comparison.
Baselines and Models.   To evaluate the effectiveness of our medical bounding box dataset for downstream radiology report generation tasks, we conducted experiments using two representative models: R2Gen and MLRG. R2Gen[36] is a Transformer-based model for medical report generation that uses relational memory and conditional layer normalization. It is a strong and widely used baseline for natural language generation tasks. MLRG[37] is a recent state-of-the-art model that leverages multi-view longitudinal data and contrastive learning. It focuses on clinical accuracy by incorporating patient-specific prior knowledge and advanced multi-modal information.
For Med-VQA tasks, we employed LLaVA-Med[38], a leading large language and vision assistant designed specifically for biomedical applications. This model excels at closed-end & open-ended medical visual question answering by effectively integrating visual information from medical images with clinical knowledge.
For the grounding tasks, we compared 5 different models: (1) the original GLIP model without fine-tuning, (2) GLIP experts fine-tuned on modality-specific subsets, (3) GLIP experts fine-tuned on 10% modality-specific subsets, (4) the original CO_DETR model without fine-tuning, and (5) CO_DETR model fine-tuned on modality-specific subsets. GLIP[29] is a vision-language model that unifies object detection and grounding by pre-training on large-scale image-text pairs, enabling open-vocabulary detection and flexible region-text alignment. Co-DETR[39] builds on the DETR architecture and incorporates language queries to perform referring expression grounding using transformers to directly link free-form text with specific image regions.
Figure 5: Performance comparison between GLIP and Med-GLIP on three medical image datasets (ClinicDB, ETIS, and Kvasir) in terms of AP@50. Med-GLIP significantly outperforms the original GLIP model across all datasets, demonstrating the effectiveness of domain adaptation for medical visual grounding tasks.
@@
Evaluation Metric.   To evaluate the precision of the grounding model GLIP, we adopted a metric called Average Precision (AP). AP is a widely used evaluation metric in object detection and grounding tasks. It measures the area under the precision-recall curve, summarizing the trade-off between precision and recall at different confidence thresholds. A higher AP indicates better overall detection or localization performance.
We adopted 2 metrics to evaluate the quality of MRG results: the natural language generation (NLG) metric and the clinical efficacy (CE) metric. The NLG metric quantifies how closely a generated report matches the reference text in terms of linguistic similarity. It includes BLEU-n, METEOR, and ROUGE-L. The CE metric evaluates whether the generated report correctly captures clinically relevant findings and their relationships. It focuses on medical accuracy rather than linguistic overlap. It includes F1 score, Precision (P), and Recall (R). Due to the MLRG model’s built-in evaluation module, this project also adopted RadGraph F1 and CheXpert F1. RadGraph F1 measures how accurately a report reproduces clinical entities and their relationships. CheXpert F1 is calculated by labeling the generated medical reports for the presence or absence of fourteen thoracic diseases and then comparing these predictions to the reference labels. For Med-VQA tasks, We employ the accuracy for closed-set questions and recall for open-set questions, being consistent with existing work like LLaVA-Med[38] for a fair comparison.

### v2

@@ Abstract
Medical image grounding aims to align natural language phrases with specific regions in medical images, serving as a foundational task for intelligent diagnosis, visual question answering (VQA), and automated report generation(MRG). However, existing research is constrained by limited modality coverage, coarse-grained annotations, and the absence of a unified, generalizable grounding framework. To address these challenges, we construct a large-scale medical grounding dataset Med-GLIP-5M comprising over 5.3 million region-level annotations across seven imaging modalities, covering diverse anatomical structures and pathological findings. The dataset supports both segmentation and grounding tasks with hierarchical region labels, ranging from organ-level boundaries to fine-grained lesions. Based on this foundation, we propose Med-GLIP, a modality-aware grounding framework trained on Med-GLIP-5M. Rather than relying on explicitly designed expert modules, Med-GLIP implicitly acquires hierarchical semantic understanding from diverse training data—enabling it to recognize multi-granularity structures, such as distinguishing lungs from pneumonia lesions. Extensive experiments demonstrate that Med-GLIP consistently outperforms state-of-the-art baselines across multiple grounding benchmarks. Furthermore, integrating its spatial outputs into downstream tasks, including medical VQA and report generation, leads to substantial performance gains. Our dataset is available at Venn2025/Med-GLIP-5M.
@@ I Introduction
With the rapid progress of medical AI, establishing precise alignment between natural language descriptions and specific regions in medical images has become a foundational step for tasks such as intelligent diagnosis, surgical navigation, and multimodal question answering [liu2021slake, zou2024medrg, chen2023medical, jiang2024moe, jiang2024med, jiang2025omniv, jiang2025hulu, jiang2019tiger, yeh2018unsupervised, chen2024r]. Medical image grounding aims to localize anatomical structures or pathological findings based on language input, enabling spatial-semantic correspondence across modalities. For example, when a radiologist states “a tumor is located in the upper right corner,” the model must accurately identify the corresponding image region. This task not only improves model interpretability but also provides critical spatial priors for downstream applications like visual question answering (VQA) and medical report generation (MRG), playing an essential role in clinical decision support[wang2025fair, chen2024r, jiang2024joint, gai2025medthink, liu2024medcot, liu2024vpl, liu2023parameter, liu2025kpl, zhang2025medu1incentivizingunifiedmedical]
However, compared to the natural image domain, medical image grounding faces several unique challenges [sun2024medical, khan2025comprehensive]. First, publicly available grounding datasets are extremely scarce, especially those with large-scale, multi-organ, and multi-modality annotations, which significantly limit progress in this field. Second, medical images are highly specialized and exhibit substantial heterogeneity across modalities such as CT, MRI, and ultrasound—in terms of spatial resolution, anatomical appearance, and contrast—making cross modal alignment inherently complex [tong2023tsnet, liu2023deep]. Moreover, many target regions, such as small lesions or vascular branches, often lack clear boundaries and exhibit high anatomical variability, placing greater demands on the model’s precision and generalization capabilities.
@@
M3D-Seg 2023 10410 studies 3D-$B$ Volumetric abnormalities
@@
To address data scarcity and semantic misalignment in medical image grounding, prior works have explored both dataset construction and cross-modal modeling. On the dataset side, SLAKE [liu2021slake] offers region-phrase annotations for chest X-rays, marking an early attempt at grounding, but it contains only 6k samples and has limited modality coverage. MedTrinity-25M [xie2024medtrinity] provides large-scale but loosely aligned image-text pairs across tasks, lacking fine-grained region-level supervision. These datasets commonly suffer from: (1) limited modality and organ diversity; (2) absence of dense region-level annotations; and (3) a focus on classification or QA tasks rather than spatial-semantic alignment. On the modeling side, recent approaches adapt natural-image cross-modal models to the medical domain (e.g., LLaVA-Med [li2023llavamed], MedKLIP [wu2023medklip], MedSAM [medsam]), often incorporating structured medical knowledge. LLaVA-Med [li2023llavamed] introduces multimodal LLMs but relies on image-caption pairs without fine-grained alignment. MedKLIP [wu2023medklip] employs region-phrase contrastive learning but remains X-ray–specific. MedSAM [medsam] incorporates structure-aware priors for zero-shot organ segmentation based on SAM [sam], yet it depends on prompts and lacks deep language grounding. OntoRay further integrates radiological ontologies to capture causal relations between terms and regions [chepelev2023ontologies]. Despite these efforts, a unified, modality-adaptive grounding framework capable of multi-scale semantic alignment and inter-modality generalization is still missing, and its impact on downstream tasks such as VQA or report generation remains underexplored.
In summary, our contributions are as follows:
@@
  • To the best of our knowledge, we introduce the largest and most diverse medical grounding dataset to date, Med-GLIP-5M, spanning 7 modalities, 30+ anatomical regions, and over 5.3 million image-text region pairs, addressing the long-standing data scarcity in this field (Table I).
  • We propose Med-GLIP, a modality-aware hierarchical expert framework that enables high-precision grounding across diverse medical imaging types by modeling fine-grained structural differences (Figure 3).
  • Through extensive experiments, we validate the effectiveness of spatial grounding in downstream tasks, showing that Med-GLIP enhances both VQA and report generation performance (Figure 1). Our work provides a unified solution across data, modeling, and application perspectives, advancing the development of generalizable medical vision-language models.
@@ II Med-GLIP-5M Construction
II-A Data Collection
We curated a multi-source medical imaging dataset for vision-language grounding. Each image was selected with a preference for multi-instance spatial annotations—either multiple masks or bounding boxes per image—thus ensuring the resulting dataset is well-suited for vision-language grounding tasks. Datasets were extracted from several platforms including but not limited to: GitHub, Grand Challenge , OpenNeuro, BraTS, SYNAPSE, and CodaLab. Selection criteria: (1) DICOM/NIfTI accessibility; (2) spatial annotations (masks/bounding boxes); (3) coverage of multiple anatomical regions. This yielded 2720979 images spanning X-ray, CT, MRI across multiple organs including abdomen(32.8%), bone(23.9%), head(17.3%), chest(7.6%) etc. The final dataset has 4.13 masks/image on average, with various modality including CT(61.84%), MRI(26.71%), Ultrasound(1.44%), X-RAY(0.22%), Dermoscopy(0.21%), Endoscopy(0.09%), Fundus(0.09%). The collected datasets span a wide range of imaging modalities, spatial resolutions, and anatomical regions, ensuring diversity across organs and clinical tasks, as shown in Figure 2.
II-B Data Quality Control
A three-tier quality control (QC) pipeline was implemented post-collection. Firstly, we discarded non-readable or broken files, sliced 3D datasets into 2D to improve data quantity, and standardized all image formats to PNG while preserving the original resolution. Secondly, we verified that each image had a corresponding and aligned segmentation mask, excluding samples that failed this check. Thirdly, we filtered inconsistent annotations, such as those with malformed structures or undefined semantics. Moreover, we removed samples where the annotated mask covered only a negligible portion of the image (empirically set as 1.5% area exclusion), as such cases do not provide meaningful spatial grounding for downstream tasks[Everingham15]. This quality assurance workflow ensures that only well-structured, clearly labeled, and spatially meaningful samples are retained for model training and evaluation.
II-C Data Preprocessing
The data preprocessing phase for Med-GLIP-5M was designed to ensure high-quality data while adhering to ethical standards and regulatory requirements. After acquisition, datasets underwent a multi-stage refinement process. Initially, data were categorized by organ type and volume, with duplicates and low-quality samples removed. Semantic label inconsistencies were resolved through standardized naming. Clean data were then augmented with geometric and intensity transformations to enhance diversity without compromising diagnostic integrity. Image formats were standardized to ensure compatibility across machine learning frameworks. A rigorous validation protocol verified the consistency between images and segmentation masks. Additionally, clinical metadata such as patient history and diagnostic outcomes were integrated where available. This structured approach resulted in a dataset that is both large-scale and high-quality, as visualized in Figure 3, providing a robust foundation for medical AI research.
@@ II-C Data Preprocessing
II-D Data Statistics
Med-GLIP-5M comprises 198 fine-grained annotation labels, which were re-organized into 38 broader anatomical categories. These 38 hierarchical classes span 6 major body regions, with a cumulative image count exceeding 11 million. This hierarchical label structure allows for flexible experimentation, such as organ-level segmentation and multi-organ detection. The dataset exhibits substantial heterogeneity with 7 distinct imaging modalities. Computed Tomography (CT) and Magnetic Resonance Imaging (MRI) are predominant, contributing approximately 4.2 million and 1.08 million images, respectively, and collectively account for over 80% of the total dataset volume. The remaining modalities, such as ultrasound, contribute important diversity.
@@ III Med-GLIP
In our work, medical object detection is reconceived as a phrase grounding task, wherein each image region that is identified corresponds to its matching medical phrase. Given a predefined set of medical concepts pertinent to a particular imaging modality, for instance, {“pneumonia”, “nodule”, “fracture”} for X-ray images, a prompt is constructed:
@@
Following methodologies similar to GLIP [li2022grounded], we can employ pre-trained language models, e.g., BERT [devlin2018bert], to encode more semantically rich prompts (e.g., “pneumonia. nodule. fracture.”), which has demonstrated empirical advantages. Within our modality-specific grounding framework, alignment scores $S_{\text{ground}}$ are computed between medical image region features $\mathbf{F}$ and the encoded word or token features $\mathbf{T}$ from the prompt. This is formally expressed as:
@@ IV-A Experiment Setup
Datasets for Region-Level Grounding.  To evaluate the grounding performance of our GLIP-experts, we selected six datasets spanning five imaging modalities: SegRap2023_ct (CT), AMOS2022_MR (MRI), sz_cxr (X-ray), CETUS2014 (Ultrasound), and CVC-ClinicDB (Endoscopy). All datasets except CVC-ClinicDB are sourced from the IMed-361M benchmark[cheng2024interactive], which provides standardized train-test splits (90%/10%). For evaluation, we use only the held-out test sets. CVC-ClinicDB lacks predefined splits and is used in its entirety. These datasets collectively cover 5,551 images with rich structural annotations across modalities.
Datasets for MRG.  We adopt two widely-used benchmarks: MIMIC-CXR[johnson2019mimic] and IU-Xray[huang2023kiut]. MIMIC-CXR includes 377,110 chest X-ray images and corresponding free-text radiology reports from 227,835 studies. It is evaluated using the MLRG framework to compute both natural language generation (NLG) metrics and concept grounding scores (e.g., CE). IU-Xray, with 7,470 image-report pairs, offers a smaller and more structured alternative, and is evaluated using the R2Gen model on standard NLG metrics.
@@
Datasets for Med-VQA.  We evaluate model generalization to vision-language reasoning via three medical VQA datasets: VQA-RAD[lau2018dataset], SLAKE[liu2021slake], and PathVQA[he2020pathvqa]. These datasets span radiological and pathological domains, include both English and Chinese annotations, and provide over 50,000 QA pairs. During evaluation, GLIP-experts support both: (1) Closed-ended multiple-choice reasoning by verifying spatial alignment between phrases and detected regions; and (2) Open-ended answer generation through grounded semantic reasoning. All evaluations follow official train-test splits for reproducibility and fair comparison.
Baselines and Models. To evaluate the effectiveness of our medical bounding box dataset for downstream radiology report generation tasks, we conducted experiments using two representative models: R2Gen and MLRG. R2Gen[chen2015generating] is a Transformer-based model for medical report generation. MLRG[liu2025enhanced] is a recent state-of-the-art model that leverages multi-view longitudinal data and contrastive learning.
For Med-VQA tasks, we employed LLaVA-Med[li2023llavamedtraininglargelanguageandvision], a leading large language and vision assistant designed specifically for biomedical applications.
For the grounding tasks, we compared 5 different models: (1) the original GLIP [li2022grounded] without fine-tuning, (2) GLIP experts fine-tuned on modality-specific subsets, (3) GLIP experts fine-tuned on 10% modality-specific subsets, (4) the original CO_DETR model[zong2023detrs] without fine-tuning, and (5) CO_DETR model fine-tuned on modality-specific subsets.
@@
Evaluation Metric. To evaluate the precision of the grounding model GLIP, we adopted a metric called Average Precision (AP). It measures the area under the precision-recall curve. A higher AP indicates better overall detection or localization performance.
We adopted 2 metrics to evaluate the quality of MRG results: the natural language generation (NLG) metric and the clinical efficacy (CE) metric. The NLG metric quantifies how closely a generated report matches the reference text in terms of linguistic similarity. It includes BLEU-n, METEOR, and ROUGE-L. The CE metric focuses on medical accuracy rather than linguistic overlap. It includes RadGraph F1 score, CheXpert F1 score, Precision (P), and Recall (R). For Med-VQA tasks, We employ the accuracy for closed-set questions and recall for open-set questions, being consistent with existing work like LLaVA-Med[li2023llavamedtraininglargelanguageandvision] for a fair comparison.